Could you please show me how to work this problem so I know how to do it.
Perimeter of a rectangle. The perimeter of a rectangular backyard is 6x + 6 yards. If the width is x yards, find the binomial that represents the length.

I need to put it in the equation forms step by step so I can understand it. Hope someone can write it for me so I understand it...
Thank you

Answer by scott8148(6628)

the perimeter is the sum of all the sides
in a rectangle , P = 2L + 2W

6x + 6 = 2L + 2x

4x + 6 = 2L

2x + 3 = L